OpenStudy (anonymous):

The formula A = 1/2 bh represents the area of a triangle where A represents the area, b is the base of the triangle and h is the height of the triangle. Solve the formula for h:

OpenStudy (anonymous):

Firstly multiply by 2 both the sides..

OpenStudy (anonymous):

2A=2b*2h ?

OpenStudy (anonymous):

2 on RHS is in denominator.

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Like: \[A = \frac{1}{2} bh\] So now multiply by 2: \[2A = 2 \times \frac{1}{2}bh\] What will you get here by solving it??

OpenStudy (anonymous):

Then divide by b both the sides to find h..

OpenStudy (anonymous):

RHS means right hand side.

OpenStudy (anonymous):

2A=bh

OpenStudy (anonymous):

A=1/2b*1/2h

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

2A = bh it is okay.. Here divide by b..

OpenStudy (anonymous):

2A/b=h

OpenStudy (anonymous):

you are dividing it by 2, divide both the sides by b..

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

and that is it?